# Random Evolutionary Dynamics in Predator-Prey Systems Yields Large, Clustered Ecosystems

### v1

@@ Abstract
We study the effect of speciation, i.e. the introduction of new species through evolution into communities, in the setting of predator-prey systems. Predator-prey dynamics is classically well modeled by Lotka-Volterra equations, also when multiple predator and prey species co-exist. The consequences of the emergence of new species in such systems are much less well understood. We find that introducing random evolving species leads to robust ecosystems in which large numbers of species coexist. Crucially, in these large ecosystems an emergent clustering of species is observed, tying functional differences to phylogenetic history.
@@ 1 Introduction
A century of quantitative work on ecosystem dynamics has produced important milestones such as the Lotka-Volterra (LV) predator-prey model [22, 43], MacArthur’s resource-consumer model [23], and May’s stability criterion [27]. The central approach in much of the quantitative modeling is to capture the time dependence of species population size via the species’ interactions with other species and the resources it consumes [41]. This often results in a system of Ordinary Differential Equations (ODEs) that describes the interaction among many species and leads to interaction or community matrices that define a community structure. The large majority of modeling work has focused on the coexistence of a fixed number of species [27]; perhaps considering the role of extinction [14]; or the invasion of new species [39], often from fixed regional pools [24]. The focus in these models often lies on the creation of communities with stable [3, 4, 6, 12, 27] and feasible equilibria [11, 16, 36]. Furthermore, they consider short timescales (in the eco-evolutionary sense), disregarding transient and evolutionary phenomena [18, 19, 30]. For a discussion of equilibrium and non-equilibrium hypotheses of biological diversity see [31].
Despite decades of research, many questions remain unanswered [4]. In recent years the role of introducing new species to the community through evolution, known as speciation, has gained traction [35, 29, 40, 32, 5]. On biological grounds, it is very natural to include speciation in ecosystem dynamics as it is common to perceive that new species come and go, yet many of the consequences of adding such dynamics remain unexplored or counterintuitive. An early attempt for a specific plankton model even showed that speciation may reduce the biodiversity in the system [34]. Note that this speciation approach is a different approach than introducing a fixed number of species with random coefficients such as [13].
Here, we show the promise of the speciation approach by showing that speciation can lead to a complex ecosystem that settles in a dynamic equilibrium, in which many species cluster into separate families of species. We explore the dynamics of speciating ecosystems that are being built up from basic principles by evolving ecosystems from just two ancestor species (predator and prey). We observe that clusters evolve out of these two ancestors and develop a phylogenetic tree of functionally different species. In our two trophic layer model system, we observe such clustering for both the predator and the prey dynamics under the condition that the variability in the speciation step is large enough.
@@
The relative importance of equilibrium and non-equilibrium dynamics has been debated for decades, see e.g. [31]. Our model shows that while its ecological and evolutionary timescales are initially well-separated in the sense that the exponential decay towards the stable state is much faster than the average time between speciation events, these timescales eventually become intertwined. As the increasing number of species leads to eigenvalues of the equilibrium with a small negative real part, we expect slow decay towards the equilibrium. Hence, speciation does not necessarily happen near a steady state, an assumption that underlies many modeling approaches such as adaptive dynamics [9]. The resulting overlap in timescales is an important feature that has a grounding in natural systems, where the ecological timescale and evolutionary timescales overlap [33, 38].
@@ 2 Model Setup
Here, $r_{i}$ is the growth rate of prey $i$ and $K$ the carrying capacity for prey $i$, which we assume to be constant for all species. The interaction between prey $i$ and predator $j$ is given by the value $P_{ij}$ which is assumed to be nonnegative.
@@ 2.1 Speciation
$(x^{*},y^{*})=\left(\frac{r}{P}\left(\frac{\delta}{\beta P},1-\frac{\delta}{\beta KP}\right)\right)=(4,1.65).$ (5)
To model the evolutionary process, we introduce a new species after a random waiting time by assuming that this new species is a perturbation of a randomly chosen ancestor species. To spawn new species from an ancestral species, choices need to be made concerning the strength of self-interaction, growth rate, species’ interaction and the sign of the species interaction. We emphasize that these model choices are crucial in determining the dynamics observed, but within a given regime of the phase space, they seem to generate similar dynamics. Notably, in [35], it is assumed that the interaction between all species is always competitive, and predator-prey and mutualism-type interactions are excluded. In other words, the interaction matrix contains only negative numbers. This is an important and restrictive model choice. However, it is also a constraint that is not easily relaxed. For instance, we find that under a broad set of conditions speciation eventually leads to blowup of the biomass in the ecosystem, if one does not constrain the sign of the interactions. We emphasize that we have no formal proof of this, but observed this dynamics for all parameter choices we could explore. As this most unbiased model choice is eventually not producing biologically realistic scenarios, we discuss it only in the Appendix A. In the main text, we limit ourselves to two trophic levels and hypothesize that imposing such a structure is a minimum requirement for maintaining a finite biomass in LV systems.
@@
Similarly, when a new predator is introduced, we add a new row to $\beta P^{T}$, resulting in a corresponding new column for $P$. We assume that the added perturbation is normally distributed with average 0 and standard deviation $\eta$. We will refer to $\eta$ as the size of the variability in the speciation step. If a newly added row to $P$ contains a negative element, it is set to 0 to keep the structure of the trophic levels. As the rows of $P^{T}$ describe how much the predators can consume, it would seem that increasing these values in a speciation step is favorable, leading to a run-off in the $P$-values. However, as $P$ and $P^{T}$ are by definition linked, a speciation event where a predator increases its positive interaction with a prey automatically increases the predation on the prey, thus counteracting the run-off.
@@
We typically take $p_{m}(0)=10^{-3}$, meaning that the average waiting time for a speciation event at the start of the simulation is $10^{3}$. Again, we emphasize that the model performance is largely insensitive to these model choices. We chose $x_{1}(0)=4$ and $y_{1}(0)=2$, hence $p_{m}(0)$ is defined with respect to a biomass of $6$, which is the order of magnitude we typically observe in the simulations as Fig. 2 in the next section will show. When a speciation event happens, the probability that species $k$ is chosen to speciate equals the fraction of its biomass to the total biomass.
Similar to the approaches in [10, 34], we assume that at a speciation event, $5\%$ of the biomass of the ancestor species $k$ transforms to the new daughter species, while the biomass of the other species does not change. We note that in the assumption that $5\%$ of the biomass of the ancestor species evolves, we implicitly assume that unsuccessful speciation events happen on a shorter timescale.
In the model, a species is declared extinct when its abundance has dropped below the extinction threshold at a new speciation event. Upon extinction of a species, its corresponding columns and rows in $P$ and $P^{T}$ are deleted (reducing the dimension of the interaction matrix $A$ from $n\times n$ to $(n-1)\times(n-1)$). Note that this step happens before the new speciation step is done. That is, an extinct species cannot speciate anymore. We typically set the extinction threshold to $10^{-3}$. A complete overview of all model settings used in our simulations (unless stated otherwise) is provided in Table 1.
@@ 3 Large, Stable and Robust Ecosystems
Figure 2: Long-time simulation of the model from sec. 2. We plot the number of species (left) and biomass (right). The total biomasses of predator and prey species appear strongly correlated and also stabilize quickly to a small slope after an initial steep growth.
For longer integration times, we only plot the state variables $S_{\mathrm{pred}}(t),S_{\mathrm{prey}}(t),B_{\mathrm{pred}}(t)$ and $B_{\mathrm{prey}}(t)$, as defined in the previous section, because figures like Fig. 1 become too cluttered at longer timescales. Fig. 2 shows the evolution of these state variables over a longer integration time of $T=3\cdot 10^{6}$. The number of predator and prey species quickly stabilizes to a ratio of approximately 1:2, but the value of the biomass keeps increasing slowly. When we repeat the simulation 100 times, with the same initial condition, we find $\langle S\rangle=81$ (rounded to the nearest integer) at $T=3\cdot 10^{6}$, but there is quite a significant spread as the histogram in Fig. 2(a) shows. Despite this spread, the histogram shows that the model consistently evolves towards an ecosystem with a high number of species. This observation and even the ratio of 1:2 is in line with real ecosystems [8]. As Fig. 1(b) indicates, there is a strong correlation between the biomasses of the predators and prey. This correlation holds not just in time, but also between different runs, as becomes clear from Fig. 2(b).
@@
To test the stability of the resulting ecosystem, we perform several tests. The first test is to stop speciation at the end of the simulation and let the system equilibrate while removing species that drop below the extinction threshold. As Fig. 3(a) shows, the ecosystem from Fig. 2 indeed equilibrates to a fixed point where 81 out of 82 species survive. Note that it takes at least up to $t=4000$ before the system becomes close to the stable state, which is significantly longer than the speciation timescale of $6661/p_{m}=$ 666 for the biomass at the start of the simulation in Fig. 3(a). Averaged over 100 simulations, 80.32 out of 80.61 species survive. This shows that the ecosystem is not dominated by species that are bound to go extinct. Furthermore, when we add a large perturbation to the stable state, for example, when we take as initial condition two times the stable state, we see that the system converges back to the stable state after a series of predator-prey cycles and no species go extinct, see Fig. 3(b). This indicates that the system is robust and is able to support large stable ecosystems.
@@ 3.1 Stationary State of the Biomass
So, we get an explicit, leading order (in $1/K$), expression for the ratio of the biomasses, which is the straight line plotted in Fig. 2(b). Heuristically, one can think of this limit as an infinite carrying capacity, i.e. $K\to\infty$, or, in other words, the absence of the logistic term. The ratio $B_{\mathrm{pred}}/B_{\mathrm{prey}}$ is identical to the ratio for the initial two-species ecosystem equilibrium, i.e. $y^{*}/x^{*}$ from Eq. (5). This approximation is justified as $1/K$ is small and $\sum_{i}x_{i}^{2}/B_{\mathrm{prey}}$ is also small because most prey species have a biomass smaller than one.
@@ 3.2 The role of variability
In ecological modeling, it is a well-known fact that variability in the ecosystem is related to niche width. In a classic work by May and MacArthur [28], it was shown that in a generalized LV model with parameters stochastically varying with variance $\sigma^{2}$, the nice width is inversely related to $\sigma^{2}$. Similar results have been found in the context of Adaptive Dynamics [1]. In our setup, the parameters do not vary stochastically within a single species but vary stochastically with deviation $\eta$ between mother and daughter species. Furthermore, as a species is defined by the rows of $A$, there is no one-dimensional niche width. However, we still expect that smaller variation, i.e. smaller $\eta$, leads to more species in the system. To test this prediction, we compute for a range of $\eta$ values, the average number of predator and prey species $\langle S\rangle$. Indeed, as Fig. 5 shows, an increase in the variability $\eta$ leads to a decrease in the number of coexisting species. Especially, the log-log plot in Fig. 4(b) shows that roughly $\langle S\rangle\sim\eta^{-1.1}$. Variability of speciation also affects the clustering ability; we will discuss this in 4.2.
@@ 3.3 Evolution in the growth rate $r$
Until now, only the values of $P$ were affected during speciation, which resulted in the construction of large stable ecosystems. We can also allow for mutation in other parameters, such as the growth rate $r_{i}$ of the prey to simulate resource competition between the prey species. In this case, we define the growth rate of a new prey species spawning from prey species $i$ as $r_{i}+\xi$ where $\xi$ is normally distributed with mean 00 and standard deviation $\eta$ (i.e. $\xi\sim\mathcal{N}(0,\eta^{2})$). Note that the same $\eta$ is used for the perturbation of $P$.
@@
We now briefly investigate the characteristics of the dynamics as was done earlier for the fixed $r$ case. The first results are shown in Fig. 6. Prey species with a higher growth rate $r_{i}$ have no direct advantage over the other prey species, as there is no explicit competition for resources. However, with a higher growth rate, the size of $x_{i}$ can become smaller before the derivative $\dot{x}_{i}$ becomes negative and the species goes extinct. Hence, over time the average value of $r$ will grow, see Fig. 11(a) in sec. 4.3. This in turn leads to more predation which explains why the biomass of the predators overtakes the biomass of the prey as shown in Fig. 5(b). On average, there are 59.58 species in the system, of which on average 58.79 remain after equilibration, see also Fig. 6(a). This average is smaller than the previous case with a fixed growth rate $r$ as there is now increased competition between the prey species. When we compare Fig. 6(b) with Fig. 2(b), we observe that the very strict correlation is broken up. However, the correlation is still well approximated by Eq. (10) when we replace $r$ with the average of $r_{i}$ over the living species at the end of the simulation, see the green line in Fig. 6(b).
@@
Summarizing, we observe that the stochastic LV model can generate realistic dynamics on the level of species and ecosystems: the model achieves a dynamic equilibrium in which speciation events are intricately coupled to species size dynamics, yet in which the overall ecosystem size is at all times realistically balanced between predators and prey. Moreover, the variability $\eta$ indeed affects the number of species in the expected manner, and the dynamic equilibrium reached is not dependent on species that without speciation would go extinct. What thus becomes possible to ask are more specific questions about the dynamics of the structure of the ecosystem. One interesting and biologically relevant question is whether a meaningful development of “clustering” emerges from the random introduction of random new species. This is the topic of the next section.

### v2

@@ Abstract
We study the effect of introducing new species through evolution into communities. We use the setting of predator-prey systems. Predator-prey dynamics is classically well modeled by Lotka-Volterra (LV) equations, also when multiple predator and prey species co-exist. We use a stochastic method to introduce new species in a two-trophic LV system. We find that introducing random evolving species leads to robust ecosystems in which large numbers of species coexist. Crucially, in these large ecosystems an emergent clustering of species is observed, tying functional differences to phylogenetic history.
@@ 1 Introduction
A century of quantitative work on ecosystem dynamics has produced important milestones such as the Lotka-Volterra (LV) predator-prey model [23, 47], MacArthur’s resource-consumer model [24], and May’s stability criterion [28]. The central approach in much of the quantitative modeling is to capture the time dependence of species population size via the species’ interactions with other species and the resources it consumes [45]. This often results in a system of Ordinary Differential Equations (ODEs) that describes the interaction among many species and leads to interaction or community matrices that define a community structure. A large part of modeling work has focused on the coexistence of a fixed number of species [28]; perhaps considering the role of extinction [14]; or the invasion of new species [43], often from fixed regional pools [25]. The focus in these models often lies on the creation of communities with stable [3, 4, 6, 12, 28] and feasible equilibria [11, 17, 40]. Furthermore, they consider short timescales (in the eco-evolutionary sense), disregarding transient and evolutionary phenomena [19, 20, 31]. For a discussion of equilibrium and non-equilibrium hypotheses of biological diversity see [32].
Despite decades of research, many questions remain unanswered [4]. In recent years the role of introducing new species to the community through evolution has gained traction [5, 30, 34, 39, 44]. For clarity of presentation, we will always call such an introduction of a new species a speciation event. In particular, we do not consider genetic variation within a species and also identify each step in adaptive diversification[10] as a speciation event in this article. On biological grounds, it is very natural to include speciation in ecosystem dynamics as it is common to perceive that new species come and go, yet many of the consequences of adding such dynamics remain unexplored or counterintuitive. An early attempt for a specific plankton model even showed that speciation may reduce the biodiversity in the system [38]. Note that this speciation approach is a different approach than introducing a fixed number of species with random coefficients such as [13].
Here, we show the promise of the speciation approach by showing that speciation can lead to a complex ecosystem that settles in a dynamic equilibrium, in which many species cluster into separate families of species. We explore the dynamics of speciating ecosystems that are being built up from basic principles by evolving ecosystems from just two ancestor species (predator and prey). We observe that clusters evolve out of these two ancestors and develop a phylogenetic tree of functionally different species. In our two trophic layer model system, we observe such clustering for both the predator and the prey dynamics under the condition that the variability in the speciation step is large enough. Note that clustering has been observed before in competitive LV models where the competition is defined as the relative position along a niche axis [33, 35, 37, e.g.].
@@
The relative importance of equilibrium and non-equilibrium dynamics has been debated for decades, see e.g. [32]. Our model shows that while its ecological and evolutionary timescales are initially well-separated in the sense that the exponential decay towards the stable state is much faster than the average time between speciation events, these timescales eventually become intertwined. As the increasing number of species leads to eigenvalues of the equilibrium with a small negative real part, we expect slow decay towards the equilibrium. Hence, speciation does not necessarily happen near a steady state. The resulting overlap in timescales is an important feature that has a grounding in natural systems, where the ecological timescale and evolutionary timescales overlap [36, 42].
@@ 2 Model Setup
Figure 1: Scheme summarizing the model setup and main outcome. 1) Population dynamics: The population dynamics are modeled with a two-trophic LV model. The prey species ($x$) has logistic growth and is consumed by predator species ($y$), which is regulated through predation rate $P_{11}$ and a conversion factor $\beta$. The predator population decreases with a linear death rate $\delta$. The equations underlying the model are further explained in sec. 2. 2) Evolution: We evolve the ecosystem by introducing new species stochastically. New species are generated by perturbing the interaction (shown) and growth rate (for prey, not shown) coefficients of a randomly selected ancestor species. The added perturbation $Q$ is drawn from a normal distribution with mean 00 and variance $\eta^{2}$. Further details on how we add evolution to the model can be found in sec. 2.1. 3) Species clustering: After many cycles of evolution and extinction, we find functional (shown) and genealogical clustering of species. The formation of clusters is discussed in sec. 4.
@@
Here, $r_{i}$ is the growth rate of prey $i$ and $K$ the carrying capacity for prey $i$, which we assume to be constant for all species. A discussion on the inclusion of the carrying capacity and its value can be found in Appendix B. The interaction between prey $i$ and predator $j$ is given by the value $P_{ij}$ which is assumed to be nonnegative.
@@
The assumption that the prey species do not compete implies that they all occupy their own niche with carrying capacity $K$, which allows us to focus on the interaction between the predators and prey. For a graphical representation of the entire model, see Fig. 1.
@@ 2.1 Speciation
$(x^{*},y^{*})=\left(\frac{\delta}{\beta P},\frac{r}{P}\left(1-\frac{\delta}{\beta KP}\right)\right)=(4,1.65).$ (5)
To model the evolutionary process, we introduce a new species after a random waiting time by assuming that this new species is a perturbation of a randomly chosen ancestor species. To spawn new species from an ancestral species, choices need to be made concerning the strength of self-interaction, growth rate, species’ interaction and the sign of the species interaction. We emphasize that these model choices are crucial in determining the dynamics observed, but within a given regime of the phase space, we see that they generate similar dynamics. Notably, in [39], it is assumed that the interaction between all species is always competitive, and predator-prey and mutualism-type interactions are excluded. In other words, the interaction matrix contains only negative numbers. This is an important and restrictive model choice. However, it is also a constraint that is not easily relaxed. For instance, we find that under a broad set of conditions, speciation eventually leads to blowup of the biomass in the ecosystem, if one does not constrain the sign of the interactions. We emphasize that we have no formal proof that the nonphysical biomass dynamics always occurs, but observed it for all parameter choices we could reasonably explore. Using the most unconstrained Lotka-Volterra model choice regarding interaction coefficients would be preferred, but as this most unconstrained choice only produces biologically unrealistic scenarios, we discuss it only in the Appendix A. In the main text, we limit ourselves to two trophic levels and hypothesize that imposing such a structure is a minimum requirement for maintaining a finite biomass in LV systems. By extension, we expect that multi-trophic levels retain the basic phenomenology of a two-trophic level system, but leave this for future work to verify.
@@
Similarly, when a new predator is introduced, we add a new row to $\beta P^{T}$, resulting in a corresponding new column for $P$. We assume that the added perturbation is normally distributed with zero average and standard deviation $\eta$. We will refer to $\eta$ as the size of the variability in the speciation step. If a newly added row to $P$ contains a negative element, it is set to zero to keep the structure of the trophic levels. As the rows of $P^{T}$ describe how much the predators can consume, it would seem that increasing these values in a speciation step is favorable, leading to a run-off in the $P$-values. However, as $P$ and $P^{T}$ are by definition linked, a speciation event where a predator increases its positive interaction with a prey automatically increases the predation on the prey, thus counteracting the run-off.
@@
We typically take $p_{m}(0)=10^{-3}$, meaning that the average waiting time for a speciation event at the start of the simulation is $10^{3}$. Again, we emphasize that the model performance is largely insensitive to these model choices. We chose $x_{1}(0)=4$ and $y_{1}(0)=2$, hence $p_{m}(0)$ is defined with respect to a biomass of $6$, which is the order of magnitude we typically observe in the simulations as Fig. 3 in the next section will show. Note that the initial condition sets $p_{m}(0)$, but does not essentially influence the simulation in any other way as the dynamics converges exponentially fast to the coexistence stable state (5). When a speciation event happens, the probability that species $k$ is chosen to speciate equals the fraction of its biomass of the total biomass.
Similar to the approaches in [9, 38], we assume that at a speciation event, a fraction $f$ of $5\%$ of the biomass of the ancestor species $k$ transforms to the new daughter species, while the biomass of the other species does not change. We note that in the assumption that $5\%$ of the biomass of the ancestor species evolves, we implicitly assume that unsuccessful speciation events happen on a shorter timescale. The outcome of the simulations is largely insensitive to the specific value of $f$ that is chosen, see Appendix B.
In the model, a species is declared extinct when its abundance has dropped below the extinction threshold at a new speciation event. Upon extinction of a species, its corresponding columns and rows in $P$ and $P^{T}$ are deleted (reducing the dimension of the interaction matrix $A$ from $n\times n$ to $(n-1)\times(n-1)$). Note that this step happens before the new speciation step is done. That is, an extinct species cannot speciate anymore. We typically set the extinction threshold to $10^{-3}$, see Appendix B for a discussion of this value. A complete overview of all model settings used in our simulations (unless stated otherwise) is provided in Table 1.
@@ 3 Large, Stable and Robust Ecosystems
Figure 3: Long-time simulation of the model from sec. 2. We plot the number of species (left) and biomass (right). The total biomass of both predator and prey species appear strongly correlated and also stabilize quickly to a small slope after an initial steep growth.
For longer integration times, we only plot the state variables $S_{\mathrm{pred}}(t),S_{\mathrm{prey}}(t),B_{\mathrm{pred}}(t)$ and $B_{\mathrm{prey}}(t)$, as defined in the previous section, because figures like Fig. 2 become too cluttered at longer timescales. Fig. 3 shows the evolution of these state variables over a longer integration time of $T=3\cdot 10^{6}$. The number of predator and prey species quickly stabilizes to a ratio of approximately 1:2, but the value of the biomass keeps increasing slowly. When we repeat the simulation 100 times, with the same initial condition, we find $\langle S\rangle=81$ (rounded to the nearest integer) at $T=3\cdot 10^{6}$, but there is a significant spread as the histogram in Fig. 4(a) shows. Despite this spread, the histogram shows that the model consistently evolves towards an ecosystem with a high number of species. This observation and even the ratio of 1:2 is in line with real ecosystems [8]. As Fig. 3(b) indicates, there is a strong correlation between the biomasses of the predators and prey. This correlation holds not just in time, but also between different runs, as becomes clear from Fig. 4(b).
@@
To test the stability of the resulting ecosystem, we perform several tests. The first test is to stop speciation at the end of the simulation and let the system equilibrate while removing species that drop below the extinction threshold. As Fig. 5(a) shows, the ecosystem from Fig. 3 indeed equilibrates to a fixed point where 81 out of 82 species survive. Numerical evaluation of the Jacobian at this fixed point also shows that all eigenvalues are negative. Furthermore, when we add a large perturbation to the stable state, for example, when we take as initial condition two times the stable state, we see that the system converges back to the stable state after a series of predator-prey cycles and no species go extinct, see Fig. 5(b). This indicates that the system is robust and is able to support large stable ecosystems. Note that it takes at least up to $t=4000$ before the system becomes close to the stable state, which is significantly longer than the speciation timescale of $6661/p_{m}=$ 666 for the biomass at the start of the simulation in Fig. 5(a). Averaged over 100 simulations, 80.32 out of 80.61 species survive. This shows that the ecosystem is not dominated by species that are bound to go extinct.
@@ 3.1 Stationary State of the Biomass
We get an explicit, leading order (in $1/K$), expression for the biomass ratio, which is the straight line plotted in Fig. 4(b). Heuristically, one can think of this limit as an infinite carrying capacity, i.e. $K\to\infty$, or, in other words, the absence of the logistic term. The ratio $B_{\mathrm{pred}}/B_{\mathrm{prey}}$ is identical to the ratio for the initial two-species ecosystem equilibrium, i.e. $y^{*}/x^{*}$ from Eq. (5). This approximation is justified as $1/K$ is small and $\sum_{i}x_{i}^{2}/B_{\mathrm{prey}}$ is also small because most prey species have a biomass smaller than one.
@@ 3.2 The role of variability
In ecological modeling, it is a well-known fact that variability in the ecosystem is related to niche width. In a classic work by May and MacArthur [29], it was shown that in a generalized LV model with parameters stochastically varying with variance $\sigma^{2}$, the niche width is inversely related to $\sigma^{2}$. Similar results have been found in [1]. In our setup, the parameters do not vary stochastically within a single species but vary stochastically with deviation $\eta$ between mother and daughter species. Furthermore, as a species is defined by the rows of $A$, there is no one-dimensional niche width. However, we still expect that smaller variation, i.e. smaller $\eta$, leads to more species in the system. To test this prediction, we compute for a range of $\eta$ values, the average number of predator and prey species $\langle S\rangle$. Indeed, as Fig. 6 shows, an increase in the variability $\eta$ leads to a decrease in the number of coexisting species. Especially, the log-log plot in Fig. 6(b) shows that roughly $\langle S\rangle\sim\eta^{-1.1}$. Variability of speciation also affects the clustering ability; we will discuss this in 4.2.
@@ 3.3 Evolution in the growth rate $r$
Until now, only the values of $P$ were affected during speciation, which resulted in the construction of large stable ecosystems. We can also allow for mutation in other parameters, such as the growth rate $r_{i}$ of the prey, to simulate resource competition between the prey species. In this case, we define the growth rate of a new prey species spawning from prey species $i$ as $r_{i}+\xi$ where $\xi$ is normally distributed with zero mean and standard deviation $\eta$ (i.e. $\xi\sim\mathcal{N}(0,\eta^{2})$). Note that the same $\eta$ is used for the perturbation of $P$.
@@
We now briefly investigate the characteristics of the dynamics as was done earlier for the fixed $r$ case. The first results are shown in Fig. 7. Prey species with a higher growth rate $r_{i}$ have no direct advantage over the other prey species, as there is no explicit competition for resources. However, with a higher growth rate, the size of $x_{i}$ can become smaller before the derivative $\dot{x}_{i}$ becomes negative and the species goes extinct. Hence, over time the average value of $r$ will grow, see Fig. 13(a) in sec. 4.3. This in turn leads to more predation which explains why the biomass of the predators overtakes the biomass of the prey as shown in Fig. 7(b). On average, there are 59.58 species in the system, of which on average 58.79 remain after equilibration, see also Fig. 8(a). This average is smaller than the previous case with a fixed growth rate $r$, as there is now increased competition between the prey species. When we compare Fig. 8(b) with Fig. 4(b), we observe that the very strict correlation is broken up. However, the correlation is still well approximated by Eq. (10) when we replace $r$ with the average of $r_{i}$ over the living species at the end of the simulation, see the green line in Fig. 8(b).
@@
Summarizing, we observe that the speciating LV model can generate realistic dynamics on the level of species and ecosystems: the model achieves a dynamic equilibrium in which speciation events are intricately coupled to species size dynamics, yet in which the overall ecosystem size is at all times realistically balanced between predators and prey. Moreover, the variability $\eta$ indeed affects the number of species in the expected manner, and the dynamic equilibrium reached is not dependent on species that without speciation would go extinct. What thus becomes possible to ask are more specific questions about the dynamics of the structure of the ecosystem. One interesting and biologically relevant question is whether a meaningful development of “clustering” emerges from the random introduction of random new species. Clustering dynamics is therefore the topic of the next section.